# Earth Observation Satellite Scheduling with Graph Neural Networks

### v1

@@ Abstract
The Earth Observation Satellite Planning (EOSP) is a difficult optimization problem with considerable practical interest. A set of requested observations must be scheduled on an agile Earth observation satellite while respecting constraints on their visibility window, as well as maneuver constraints that impose varying delays between successive observations. In addition, the problem is largely oversubscribed: there are much more candidate observations than what can possibly be achieved. Therefore, one must select the set of observations that will be performed while maximizing their weighted cumulative benefit, and propose a feasible schedule for these observations. As previous work mostly focused on heuristic and iterative search algorithms, this paper presents a new technique for selecting and scheduling observations based on Graph Neural Networks (GNNs) and Deep Reinforcement Learning (DRL). GNNs are used to extract relevant information from the graphs representing instances of the EOSP, and DRL drives the search for optimal schedules. Our simulations show that it is able to learn on small problem instances and generalize to larger real-world instances, with very competitive performance compared to traditional approaches.
@@ 1 Introduction
An Earth observation satellite (EOS) must acquire photographs of various locations on the surface of Earth to satisfy user requests. An agile EOS has degrees of freedom allowing it to target locations that are not exactly at its vertical in an earth-bound referential (“nadir”). The satellite we consider is in low orbit; as a consequence, each observation is available in a visible time window (VTW) that is significantly larger than its acquisition duration. Maneuvering the satellite between two observations consists of modifying its pitch, yawn and roll triple—also called its attitude—and thus implies delays that depend on the start and end observation, as well as on the time of the transition [18]. In addition, an agile EOS is typically oversubscribed: there is more observations to be performed that can possibly be achieved in the given time of operation. As different acquisitions may be associated with different priorities or utilities, the Earth observation satellite planning problem (EOSP) consists in selecting a set of acquisitions that maximize their weighted cumulative values, and designing a schedule for acquiring these observations while respecting the operational constraints.
@@
The most complex instances of the EOSP involve several satellites orbiting Earth over multiple orbits, and dependencies between targets [22]. For instance, an acquisition may consist of several pictures of the same earth-bound location to be taken by different satellites, and/or in different time-windows [19]. In this paper, we limit our study to single-satellite, single-orbit, and single-shot problems, where there is only one satellite to control over a single orbit, and each acquisition is made of a single picture to be taken in its given VTW. Nevertheless, the problem is NP-complete [13], and there is no practical solution to compute optimal schedules for problems of realistic size which contain a few thousands of candidate acquisitions, thus focusing previous work towards approximate, heuristic and random search algorithms [22]. Variants of the greedy randomized adaptive search procedure (GRASP) [6] are commonly deployed in practical applications. At the same time, the field of combinatorial optimization is currently the subject of an accrued interest from researchers in deep learning. In particular, Deep Reinforcement Learning (DRL) offers a framework to learn heuristics for NP-complete problems that has been successfully applied to a wide range of problems [24]. Following this trend, we build on previous work using a state-of-the-art combination of graph neural networks (GNN) and DRL. This approach code-named Wheatley was developed to address Job Shop Scheduling Problems with duration uncertainty [11]. Here we adapt it to the EOSP and prove its efficiency in solving deterministic but largely over-subscribed problems. Our simulation results show that we outperform the currently deployed techniques.
@@
Our main contributions are threefold: (1) propose a graph search representation of the problem without discretizing time; (2) use deep reinforcement learning to solve the problem; (3) use directly graph representation as observations thanks to graph neural networks. The main outcomes are: (1) very competitive results compared to baselines; (2) good generalization abilities allowing to train on small instances an solve efficiently large instances.
@@
The paper is organized as follows. First we give a quick survey of previous work using deep learning to solve the EOSP. Then, in Section 3 ,we introduce the problem and discuss various representations used in this work. Section 4 is dedicated to the description of the machine learning architecture used for optimization. We provide simulation results on large size real-world instances of the problem in Section 5. We conclude and discuss further research directions in Section 6.
@@ 2 Related Work
The EOSP has been subject to a large body of research, from communities as varied as aerospace and engineering, operational research, computer science, remote sensing and multipliscinary sciences. We refer the reader to [22] for a survey of non-machine learning approaches to the problem, and we focus our attention on DRL based approaches to the EOSP. Note that we address the time-dependent EOSP, where the duration of a transition between two observations varies with time. This contrasts with most of previous literature that assumes constant, time-independent transition duration.
Peng et al. [14] address a slightly different problem where observations are scheduled on board. A LSTM-based encoding network is used to extract features, and a classification network is used to make the a decision. Dalin et al. [4] solve multi-satellite instances by modeling them as a Multi-Agent Markov Decision Processes, then use a DRL actor-critic architecture. The actor is decentralized, each satellite using a relatively shallow network to select its action. The critic is centralized and implemented as a large recurrent network taking input from all satellites. Hermann et al. [9] also address the multi-satellite problem: a policy is trained in a single satellite environment on a fixed number of imaging targets, and then deployed in a multi-satellite scenario where each spacecraft has its own list of imaging targets. Local policies are learned using a combination of Monte Carlo Tree Search (MCTS) to produce trajectories, and supervised learning to learn Q-values using the trajectories produced by MCTS as training examples. Finally, Chun et al. [3] present a very similar approach; the main difference is that the transitions durations approximated during training phase instead of being precisely computed on discrete dates before training.
@@ 3 Problem Representation
An instance of the EOSP is defined by a set of candidate observations, or acquisitions, and a time-horizon $\tau$ (in this work, the duration of an orbit). Each observation $i$ is associated with its fixed duration $d_{i}$ and its VTW $e_{i};l_{i}]$ such that $l_{i}\leq\tau$. The transition duration between two acquisitions $i$ and $j$ is a function $\Delta_{i,j}(t_{i})$ of the starting time $t_{i}$ of the first observation. A schedule is a sequence of observations with associated starting time. It is represented as a single mapping that associates with each candidate observation $i$ its starting time $t_{i}$, such that $t_{i}=-1$ for all observations $i$ not included in the schedule. A schedule is feasible if: (i) each scheduled observation starts and ends within its VTW: $t_{i}\in[e_{i};l_{i}-d_{i}]$; (ii) the time-gap between two successive observations is greater or equal to the transition delay: $\forall i,j$ such that $t_{i}\neq-1$, $t_{j}\neq-1$ and $t_{i}\leq t_{j}$ then $t_{j}-t_{i}\geq d_{i}+\Delta_{i,j}(t_{i})$. Each observation is associated with a utility value $u_{i}$ in $\mathbb{R}^{+}$. The goal is then to find a feasible schedule that maximizes the cumulative utility of the observations it includes: $\mathit{maximize}\left(\sum_{i:t_{i}\neq-1}u_{i}\right)$.
Figure 1: Discrete graph for 4 candidate acquisitions. Corresp. continuous graph is a 4-node clique.
@@ 3.1 Classical Approach : Time Discretization
In the EOSP, the start time $t_{i}$ of each observation is a continuous value, and the transition durations $\Delta_{i,j}(t_{i})$ are continuous functions of continuous variables. Therefore, the EOSP is not a completely discrete problem. However, it is often re-framed as such, either by making assumptions on the start time of transitions (for instance, every transition starts as soon as it is available), or by crudely discretizing the time variable.
@@
In this work, the problem is provided under the form a very large time-discretized graph that represents the EOSP as a sequential decision problem. In this graph—later called the "discrete graph"—every candidate acquisition is represented once for each possible (discrete) start time. A node is typically denoted $(i,t_{i})$ and represents starting observation $i$ at time $t_{i}\in[e_{i},l_{i}-d_{i}]$. An edge between two nodes is present if the end observation is a possible immediate successor of the start observation. The acquisition and transition durations are accounted for while defining the edges: an edge between $(i,t_{i})$ and $(j,t_{j})$ is such that $t_{j}$ is the smallest discrete time satisfying $t_{i}+d_{i}+\Delta_{ij}(t_{i})\leq t_{j}$ and $t_{j}\geq e_{j}$. Note that there is an (implicit) mutual exclusion between two nodes $(i,t_{i})$ and $(i,t_{i}^{\prime})$, $t_{i}\neq t_{i}^{\prime}$, to indicate that observation $i$ must not be performed twice. Finally, some edges are pruned using considerations of optimality: if an observation $k$ can be inserted between $(i,t_{i})$ and $(j,t_{j})$ without breaking the constraints of the problem (that is, the transition delays), then the edge between $(i,t_{i})$ and $(j,t_{j})$ is removed. In this case, every path between the two nodes must go through one node $(k,t_{k})$. The reasoning is that every optimal solution that includes both $(i,t_{i})$ and $(j,t_{j})$ would also include observation $k$. Therefore, $(j,t_{j})$ should not be an immediate successor of $(i,t_{i})$. Figure 1 provides an example of discrete graph for a problem containing 4 candidate acquisitions.
The discrete graph is convenient for a typical state-space approach such as using an implicit enumeration algorithm (Dikstra, A*). A schedule can be built by starting from a node representing the origin of the graph and following the edges to add observations in chronological order (the earliest are added to the scheduled before the latest). Let $m$ denote the last node scheduled. When a new node $n=(i,t_{i})$ is scheduled, the graph is updated in the following way: (1) all edges outgoing from node $m$ are removed, except the one leading to the newly scheduled node $n$; (2) all other nodes candidate for observation $i$ (nodes $(i,t^{\prime}_{i})$ with $t^{\prime}i\neq t_{i}$) are deleted; (3) all nodes that have become unreachable from the origin node are removed.
@@ 3.2 Our Approach : Continuous-Time Graph
Although convenient for state-space approaches, the discrete graph has the drawback of quickly becoming huge as the number of candidate acquisition grows, making it unsuitable as an input to a GNN. For this reason, we derive from the discrete graph a (much) more compact graph that we call the "continuous graph". In this graph, each candidate observation is represented exactly once, with no mention of its exact starting time. A node is simply defined by a candidate observation $i$. An edge $(i,j)$ is present in the continuous graph if and only if there is an edge $((i,t_{i}),(j,t_{j}))$ in the discrete graph. Note that this may lead to two nodes pointing at each other, if the corresponding observations may be performed in any order.
At each iteration of our algorithm, the agent is fed with the compact continuous graph and selects the next acquisition to be inserted in the schedule. To track which candidate observations are disallowed by this decision, and what are the possible next actions, we implement this decision in the discrete graph by assuming the selected observation is started at its earliest possible (discrete) time. The discrete graph is updated as described above, and these changes are reflected in the continuous graph. The resulting continuous graph is fed to the GNN at the next iteration. At each step, the set of candidate acquisitions that can be added to the current schedule is the set of immediate successors of the last node scheduled in the discrete graph. The learning algorithm is responsible for selecting one particular acquisition among them, using the procedure detailed later.
@@
As stated above, in this work, we rely on the discrete graph. We use it both for building the continuous graph, and as a simulator of the built strategy. Pre-building this discrete graph could be avoided by building directly the continuous graph from the problem data, and using a satellite simulator to get the attitudes and transition times on-the-fly, without significant change in our approach. At the time of writing, the discrete graph is precomputed by calling a closed-source proprietary satellite simulator, and there is no simple legal way to switch to the “on-the-fly” approach.
3.3 Sequential Decision Model
Our solution technique is based on representing the process of building an optimal schedule for an instance of EOSP as a reinforcement learning problem, then using DRL algorithms to solve it. Reinforcement learning is concerned with learning the solution of a Markov Decision Process (MDP), which is a discrete-time sequential decision model. An MDP is defined as a tuple $(S,A,T,R)$ where $S$ is a state space, $A$ the action space, $T$ the transition matrix, and $R$ the reward function [15]. The definition of these elements flows directly from the representation of the EOSP as a discrete graph:
  • A state $s\in S$ is a discrete EOSP graph as defined before. It can be either the initial graph where no task has been selected yet, or an intermediate graph where some nodes and edges have been selected to account for the fact that the beginning of the schedule has been decided;
  • Given a state $s$, the set of available actions $a$ is the set of all possible successors of the last selected node in $s$ (thus leading to chronological insertions);
  • MDPs naturally handle uncertainty in the problem. I the general case, it is represented in the transition matrix: $T(s,a,s^{\prime})=\Pr(s(t+1)=s^{\prime}\mid s(t)=s,a(t)=a)$. However, our formulation of the EOSP is deterministic, therefore, the MDP we derive contains no uncertainty. Given an initial state $s$ (discrete graph) and selected action $a$ (the next observation to add to the schedule), the transition matrix gives probability 1 to the state $s^{\prime}$ representing the discrete graph after the addition of $a$ to the schedule, following the update procedure described in the previous section.
  • As every inserted observation is feasible by definition, we use as immediate reward the utility value associated with the selected observation ie $R(s,a,s^{\prime})=u_{a}$. The aim is to maximize the undiscounted sum of immediate rewards.
While these components define a fully observable MDP, we do not provide a perfect description of the state $s$ to the deep network. As explained above, we do not feed the huge discrete graph to the neural network. Instead, we use a continuous graph that is not a perfectly accurate representation of the problem, because transition duration are not reported exactly (see section 4.2). Therefore, the learner solves partially-observable MDP (POMDP) [12], where partial observability concerns only the transition durations, and thus plays a minor role.

### v2

# Improved Earth Observation Satellite Scheduling with Graph Neural Networks and Monte Carlo Tree Search

@@ Abstract
Earth Observation Satellite Planning (EOSP) is a difficult optimization problem with considerable practical interest. A set of requested observations must be scheduled on an agile Earth observation satellite while respecting constraints on their visibility window, as well as maneuver constraints that impose varying delays between successive observations. In addition, the problem is largely oversubscribed: there are much more candidate observations than can possibly be achieved. Therefore, one must select the set of observations that will be performed while maximizing their cumulative benefit and propose a feasible schedule for these observations. As previous work mostly focused on heuristic and iterative search algorithms, this paper presents a new technique for selecting and scheduling observations based on Graph Neural Networks (GNNs) and Deep Reinforcement Learning (DRL). GNNs are used to extract relevant information from the graphs representing instances of the EOSP, and DRL drives the search for optimal schedules. A post-learning search step based on Monte Carlo Tree Search (MCTS) is added that is able to find even better solutions. Experiments show that it is able to learn on small problem instances and generalize to larger real-world instances, with very competitive performance compared to traditional approaches.
@@ 1 Introduction
An Earth observation satellite (EOS) must acquire photographs of various locations on the surface of Earth to satisfy user requests. An agile EOS has degrees of freedom allowing it to target locations that are not exactly at its vertical in an earth-bound referential (“nadir”). The satellite we consider is in low orbit; as a consequence, each observation is available in a visibility time window (VTW) that is significantly larger than its acquisition duration. Maneuvering the satellite between two observations consists of modifying its pitch, yaw and roll angles triple—also called its attitude—and thus implies delays that depend on the start and end observation targets as well as on the maneuver start date (time_dependent_eosp). In addition, an agile EOS is typically oversubscribed: there are more observations to be performed that can possibly be achieved in the given operation temporal horizon. As different acquisitions may be associated with different priorities or utilities, the Earth observation satellite planning problem (EOSP) consists in selecting a set of acquisitions that maximize their weighted cumulative values and designing a schedule for acquiring these observations while respecting the operational constraints.
@@
The most complex instances of the EOSP involve several satellites orbiting the Earth over multiple orbits, and dependencies between targets (EOSP_20_years). For instance, an acquisition may consist of several pictures of the same earth-bound location to be taken by different satellites, and/or in different time-windows (squillaci2023scheduling). In this paper, we limit our study to single-satellite, single-orbit, and single-shot problems, where there is only one satellite to control over a single orbit, and each acquisition is made of a single picture to be taken in its given VTW. Nevertheless, the problem is NP-complete (lemaitre2002selecting), and there is no practical solution to compute optimal schedules for problems of realistic size which contain a few thousand candidate acquisitions, thus focusing previous work towards approximate, heuristic and random search algorithms (EOSP_20_years). Variants of the greedy randomized adaptive search procedure (GRASP) (grasp) are commonly deployed in practical applications. In (Pralet23), the author considers a single satellite scheduling problem with time-dependent maneuvers and aims at minimizing the tardiness associated with a set of observations that must be performed. The solving approach combines Dynamic Programming and Large Neighborhood Search techniques. The problem we consider here is different as decisions must not only be made on the observations scheduling order but also on their presence in the final scheduling.
At the same time, the field of combinatorial optimization is currently the subject of an accrued interest from researchers in deep learning. In particular, Deep Reinforcement Learning (DRL) offers a framework for learning heuristics for NP-complete problems that has been successfully applied to a wide range of problems (nco). Following this trend, we build on previous work using a state-of-the-art combination of graph neural networks (GNN) and DRL. This approach code-named Wheatley was developed to address Job Shop Scheduling Problems with duration uncertainty (wheatley-jssp). Here we adapt it to the EOSP and prove its efficiency in solving deterministic but largely over-subscribed problems. Our simulation results show that we outperform the currently deployed techniques.
@@
Our main contributions are as follows: (1) propose a graph search representation of the problem without time discretization; (2) use deep reinforcement learning to solve the problem; (3) use directly the problem graph representation as observations within the graph neural network; (4) develop a post-training search phase based on MCTS that substantially improves the results. The main outcomes are: (1) very competitive results compared to baselines; (2) good generalization abilities allowing to train on small instances and solve efficiently large instances.
@@
The paper is organized as follows. First, we give a quick survey of related work based on deep learning approaches to solve the EOSP. Then, in Section 3, we introduce the problem and discuss various representations used in this work. Section 4 is dedicated to the description of the machine learning architecture used for optimization and post-training search. We provide simulation results on large size real-world instances of the problem in Section 5. We finally conclude and discuss further research directions.
@@ 2 Related Work
The EOSP has been subject to a large body of research, from communities as varied as aerospace and engineering, operational research, computer science, remote sensing and multidisciplinary sciences. We refer the reader to (EOSP_20_years) for a survey of non-machine learning approaches to the problem, and we focus our attention on DRL based approaches to the EOSP. Note that we address the time-dependent EOSP, where the duration of a transition between two observations varies with time. This contrasts with most of previous literature that assumes constant, time-independent transition duration.
Peng et al. (EOSP_LSTM) address a slightly different problem where observations are scheduled on board. A LSTM-based encoding network is used to extract features, and a classification network is used to make a decision. Dalin et al. (Multiagent_DRL) solve multi-satellite instances by modeling them as a Multi-Agent Markov Decision Processes, then use a DRL actor-critic architecture. The actor is decentralized, each satellite using a relatively shallow network to select its action. The critic is centralized and implemented as a large recurrent network taking input from all satellites. Hermann et al. (RL_Agile) also address the multi-satellite problem: a policy is trained in a single satellite environment on a fixed number of imaging targets, and then deployed in a multi-satellite scenario where each spacecraft has its own list of imaging targets. Local policies are learned using a combination of Monte Carlo Tree Search (MCTS) to produce trajectories, and supervised learning to learn Q-values using the trajectories produced by MCTS as training examples. Finally, Chun et al. (math11194059) present a very similar approach; the main difference is that the transition durations are approximated during the training phase, whereas in our approach they are precisely computed based on discrete date values before training.
@@ 3 Problem Representation
An instance of the EOSP is defined by a set of candidate observations $\mathcal{O}$, or acquisition requests, and a time-horizon $\tau$ (in this work, the duration of an orbit). Each observation $i\in\mathcal{O}$ is associated with its fixed duration $d_{i}$ and its VTW $e_{i};l_{i}]$ such that $l_{i}\leq\tau$.
The transition duration between two acquisitions $i$ and $j$ is a function $\Delta_{i,j}(t_{i})$ of the starting time $t_{i}$ of the first observation  [^1] Strictly speaking, the function depends on the starting time of the maneuver, which is equal to $t_{i}+d_{i}$, where $d_{i}$ is a deterministic duration.. A schedule $\sigma$ is a sequence of selected observations with associated starting time. It is represented by a single mapping that associates with each candidate observation $i\in\mathcal{O}$ its starting time $t_{i}^{\sigma}$, such that $t_{i}^{\sigma}=-1$ for all observations $i$ not selected in schedule $\sigma$. When there is no ambiguity on the schedule $\sigma$ considered, we write $t_{i}$ instead of $t_{i}^{\sigma}$. A schedule $\sigma$ is feasible if: (i) each scheduled observation starts and ends within its VTW: $\forall i\in\mathcal{O}$ such that $t_{i}\neq-1$, $t_{i}\in W_{i}$ where $W_{i}=[e_{i};l_{i}-d_{i}]$; (ii) the time gap between two successive selected observations is greater or equal to the transition delay: $\forall i,j\in\mathcal{O}^{2}$ such that $t_{i}\neq-1$, $t_{j}\neq-1$ and $t_{i}\leq t_{j}$, $t_{j}-t_{i}\geq d_{i}+\Delta_{i,j}(t_{i})$.
Each observation is associated with a utility value $u_{i}$ in $\mathbb{R}^{+}$. The goal is then to find a feasible schedule that maximizes the cumulative utility of the observations it includes: $\mathit{maximize}\left(\sum_{i:t_{i}\neq-1}u_{i}\right)$.
@@ 3.1 Classical Approach : Time Discretization
In the EOSP, the start time $t_{i}$ of each observation has a continuous domain ($W_{i}$), and the transition durations $\Delta_{i,j}(t_{i})$ are continuous functions of continuous variables. Therefore, the EOSP is not a pure discrete problem. However, it is often re-framed as such, either by making assumptions on the start time of transitions (for instance, every transition starts as soon as possible), or by crudely discretizing the time variable domains.
Discrete graph.
@@
In this work, the problem input is provided under the form a very large time-discretized graph, later called the “discrete graph”, and denoted $\mathcal{G}^{D}$. More precisely, each visibility window $W_{i}$ associated with observation $i$ is discretized into a set of candidate starting dates denoted $W_{i}^{D}$. In graph $\mathcal{G}^{D}$, every candidate acquisition $i$ is represented $\mathit{card}(W_{i}^{D})$ times. Formally, for each observation $i$ and each possible discretized starting time $t_{i}\in W_{i}^{D}$, $\mathcal{G}^{D}$ contains a node $(i,t_{i})$ representing the fact that observation $i$ start date is equal to $t_{i}$. An arc between two nodes associated with observations $i$ and $j$ is present if the end observation $j$ is a possible immediate successor of the start observation $i$. The acquisition and transition durations are accounted for while defining the arcs: an arc between $(i,t_{i})$ and $(j,t_{j})$ is such that $t_{j}$ is the smallest discrete time in $W_{j}^{D}$ satisfying $t_{i}+d_{i}+\Delta_{ij}(t_{i})\leq t_{j}$. Note that there is an (implicit) mutual exclusion between two nodes $(i,t_{i})$ and $(i,t_{i}^{\prime})$, $t_{i}\neq t_{i}^{\prime}$, to indicate that observation $i$ must not be performed twice. A virtual source node $(0,t_{0})$ is added to represent the temporal horizon starting date.
Finally, some arcs are pruned using considerations of optimality: if an observation $k$ can be inserted between $(i,t_{i})$ and $(j,t_{j})$ without breaking the constraints of the problem (that is, the transition delays), then the arc between $(i,t_{i})$ and $(j,t_{j})$ is removed. In this case, every path between the two nodes must go through one node $(k,t_{k})$. The reasoning is that every optimal solution that includes both $(i,t_{i})$ and $(j,t_{j})$ would also include observation $k$. Therefore, $(j,t_{j})$ should not be an immediate successor of $(i,t_{i})$. Figure 1 provides an example of a discrete graph for a problem containing 4 candidate acquisitions.
Figure 1: Discrete graph for 4 candidate acquisitions. Corresponding continuous graph is a 4-node clique.
Schedule and graph update.
With such a discrete graph, a schedule can be represented as path starting from the source node. Formally, a schedule is a list of nodes $((0,t_{0}),(k,t_{k}),\ldots(n,t_{n}))$ in which $(0,t_{0})$ is the source node and successive pairs of nodes in the path are arcs. Nodes that are not part of the path are considered as not selected. When building sequentially a schedule, the discrete graph can be simplified accordingly, as detailed as follows. Let $m$ denote the last node in a schedule $\sigma$. When adding a new node $n=(i,t_{i})$ to $\sigma$, the graph is updated in the following way: (1) all arcs outgoing from node $m$ are removed, except the one leading to the newly scheduled node $n$; (2) all other nodes candidate for observation $i$ (nodes $(i,t^{\prime}_{i})$ with $t^{\prime}i\neq t_{i}$) are deleted; (3) all nodes that are unreachable from $n$ are removed.
The discrete graph is convenient for a typical state-space approach such as using an implicit enumeration algorithm (Dijkstra, A*) or Dynamic Programming. Our solution technique is based on representing the process of building an optimal schedule for an instance of the EOSP as a reinforcement learning problem.
3.2 Sequential Decision Model
Reinforcement learning is concerned with learning the solution of a Markov Decision Process (MDP), which is a discrete-time sequential decision model. An MDP is defined as a tuple $(S,A,T,R)$ where $S$ is the state space, $A$ the action space, $T$ the transition matrix, and $R$ the reward function (puterman2014markov). The definition of these elements flows directly from the discrete graph representation:
  • A state $s\in S$ is a discrete EOSP graph as defined before along with a schedule $\sigma$. It can be either the initial discrete graph for which the schedule is empty, or an updated graph along with a non-empty schedule;
  • Given a state $s$ and its associated schedule $\sigma$, the set of available actions $a$ is the set of all possible successors of the last node in $\sigma$;
  • MDPs naturally handle uncertainty in the problem. In the general case, it is represented in the transition matrix: $T(s,a,s^{\prime})=\Pr(s(t+1)=s^{\prime}\mid s(t)=s,a(t)=a)$. However, our formulation of the EOSP is deterministic, therefore the MDP contains no uncertainty  [^2] Note that non-deterministic MDPs could be considered in the EOSP to account for the cloud cover uncertainty. Such an aspect is out of scope in this work.. Given an initial state $s$ (discrete graph with schedule) and a selected action $a$ (the next observation to add to the schedule), the transition matrix gives probability 1 to the state $s^{\prime}$ representing the discrete graph after the addition of $a$ to the schedule, following the update procedure described previously.
  • As every inserted observation is feasible by definition, we use as an immediate reward the utility value associated with the selected observation i.e. $R(s,a,s^{\prime})=u_{a}$. The aim is to maximize the undiscounted sum of immediate rewards.
These components define a fully observable MDP on which RL approaches can be based: the agent learns how to choose the next action in each state (policy) and the environment (or simulator) is responsible for providing a reward associated with each selected action and updating the state accordingly. Although convenient for state-space approaches, the discrete graph has the drawback of quickly becoming huge as the number of candidate acquisition grows, making it unsuitable as an input to a GNN-based agent. For this reason, we derive from the discrete graph $\mathcal{G}^{D}$ a (much) more compact graph $\mathcal{G}^{C}$ that we call the “continuous graph” and use it as an input for the agent.
@@ 3.3 Our Approach : Continuous-Time Graph
The continuous graph $\mathcal{G}^{C}$ is built upon the discrete graph $\mathcal{G}^{D}$ as follows. In $\mathcal{G}^{C}$, a node is simply defined by a candidate observation $i$, with no mention of its exact starting time. An arc $(i,j)$ is present in $\mathcal{G}^{C}$ if and only if there is an arc $((i,t_{i}),(j,t_{j}))$ in $\mathcal{G}^{D}$. Note that this may lead to two-nodes-cycles if the corresponding observations may be performed in any order. In $\mathcal{G}^{C}$, the set of every cycle-free path is a super set of the possible schedules: after selecting some acquisitions, VTW for later acquisitions may shrink to empty time windows, making these later acquisitions impossible to select.
The continuous graph $\mathcal{G}^{C}$ cannot be used only by itself in the RL approach, as it does not contain the precise transition duration information. However, it is compact enough to be used as an input for a GNN-based agent. Therefore, our RL approach uses both graph representations. As illustrated on Figure 2, the agent is fed with the compact continuous graph $\mathcal{G}^{C}$ and selects the next acquisition $i$ to be inserted in the schedule. To track which candidate observations are the possible next actions, the simulator uses the discrete graph $\mathcal{G}^{D}$ and assumes that observation $i$ is started at its earliest possible (discrete) time. The discrete graph $\mathcal{G}^{D}$ is updated with this decision as described previously, and these changes are then reflected in the continuous graph $\mathcal{G}^{C}$. The resulting continuous graph is fed to the GNN-based agent at the next iteration. The set of candidate acquisitions that can be added to the current schedule is the set of immediate successors of the last node scheduled in the discrete graph $\mathcal{G}^{D}$. Note that node features of the continuous graph contain information about transition durations (see Section 4.2). In terms of the associated MDP, the learner solves partially-observable MDP (POMDP) (kaelbling:aij98), where partial observability concerns only the transition durations, and thus plays a minor role.
@@
In the approach presented here, the discrete graph is used both for building the continuous graph and as a simulator of the learned strategy. Having a satellite simulator capable of getting the attitudes and transition times on-the-fly would remove the discrete graph building requirement, without significant change in our approach. At the time of writing, the discrete graph is precomputed by calling a closed-source proprietary satellite simulator, and there is no simple legal way to switch to the “on-the-fly” approach.